# Leveraging Segment Anything Model for Source-Free Domain Adaptation via Dual Feature Guided Auto-Prompting

### v2

@@ 1 Introduction
In general, there are two ways of exploiting SAM in the SFDA scenario. One way is to finetune SAM (usually by injecting adapters), with the algorithms in, e.g., SAMed [18], MA-SAM [19] and Med-SA [20]. However, in SFDA where the ground truth is unavailable, the performance of finetuning SAM highly depends on the quality of pseudo-labels. Bad pseudo-labels cannot yield satisfactory results (see Section 5.3.3). Another manner of leveraging SAM for SFDA is generating superior segmentation masks with the counterparts of SAM in the medical domain, e.g., MedSAM [21] and SAM-Med3D [22], to improve SFDA results, where the bottleneck is auto-generated valid prompts. Among the prompting modes of the medical domain SAM, the bounding box prompt can clearly specify the segmentation target without ambiguity and yields better segmentation accuracy compared to the point prompt[21, 22]. Thus, we aim to offer bounding box prompts to SAM, such as MedSAM[21], for enhancing domain adaptation. Existing auto-prompting methods require supervised fine-tuning of SAM, which falls outside the scope of the SFDA paradigm[23, 24]. Naively using the predictions of the model resulting from existing SFDA methods to obtain box prompts for SAM shows limited performance gains (see the rows “ProtoContra w/ MedSAM” in Table 123). Therefore, a better method of automatically providing the bounding box prompt needs to be designed. This paper focuses on using MedSAM, a renowned medical domain SAM with box prompt.
@@ 4.1 Preliminaries
Denote the source domain by $\mathcal{D}^{s}=(\mathcal{X}^{s},\mathcal{Y}^{s})$ and target domain by $\mathcal{D}^{t}=(\mathcal{X}^{t},\mathcal{Y}^{t})$. In the SFDA problem, a source model $F^{s}:\mathcal{X}^{s}\rightarrow\mathcal{Y}^{s}$ is trained with a source domain dataset $\{x^{s}_{i},y^{s}_{i}\}_{i=1}^{n_{s}}$, where $(x^{s}_{i},y^{s}_{i})\in(\mathcal{X}^{s},\mathcal{Y}^{s})$. In this paper, we consider the medical image segmentation task, so $F^{s}$ is usually trained with a supervision loss like cross-entropy loss or Dice loss. $F^{s}$ consists of two parts: a feature extractor $G^{s}$ and a classifier $H^{s}$, i.e., $F^{s}(x)=H^{s}(G^{s}(x))$. $G^{s}$ maps each pixel $i\!\in\!\{1,\!\cdots\!,h\!\times\!w\}$ in an image to the feature $f_{i}\!\in\!\mathbb{R}^{d_{F}}$, and $H^{s}$ projects pixel feature into the semantic label space with $K$ classes. Here $h\times w$ is the image size and $d_{F}$ is the dimension of the extracted feature in model $F$. The source model is provided to the target client side. Also, an unlabeled dataset $\{x^{t}_{i}\}_{i=1}^{n_{t}}$ from the target domain $\mathcal{D}_{t}$ is given, where $x^{t}_{i}\in\mathcal{X}^{t}$. SFDA aims to learn a target model $F^{t}:\mathcal{X}^{t}\rightarrow\mathcal{Y}^{t}$ with only the source model $F^{s}$ and the target unannotated dataset $\{x^{t}_{i}\}_{i=1}^{n_{t}}$.
@@ 5.2 Comparison with State-of-the-Art Methods
To validate the effectiveness of our proposed DFG, we compared it with five recent state-of-the-art (SOTA) SFDA for medical image segmentation methods: DPL [10], AdaMI [15], UPL [12], FVP [14], and ProtoContra [16]. DPL [10] denoised pseudo-labels via pixel uncertainty estimation and prototype estimation. AdaMI [15] minimizes the entropy of the model predictions with a class-ratio regularizer. UPL [12] obtain pseudo labels and their uncertainty via diversified predictions. FVP [14] learns a visual prompt parameterized with some low-frequency parameters in the frequency space. ProtoContra [16] aligns the target features with class prototypes and uses contrastive learning to train the pixels with unreliable predictions. For a fair comparison, we reimplement DPL, AdaMI, UPL, and ProtoContra based on their papers and their released codes. We utilize the same segmentation network, data augmentations, and type of optimizers. The batch sizes are tuned to choose the optimal values, and all the other hyperparameters are unchanged. For FVP, we directly reference the results from their paper as the codes are not released.
@@ 5.3.3 Comparison with finetuning Segment Anything Model
Figure 7: Examples of failure cases of box prompt search. From left to right: Target images; Ground truth; The ground truth box prompt (red box), and the procedure of box prompt search (blue boxes), with the boxes changing from small to large (For clarity, only the first, the last, and part of the other prompts are drawn); The number of pixels of changed MedSAM predictions when the box prompt is switched from the last to the current, with the horizontal axis being the index of the current prompt.
@@ 5.3.6 Analysis of efficiency and parameters
In terms of the number of parameters, our DFG freezes the MedSAM and trains only the target model without introducing additional trainable networks for adaptation. Namely, the trainable parameters are limited to those of the target segmentation network, which are the same as or fewer than previous SFDA methods, with the exception of FVP, which solely trains a visual prompt. Despite this, our method surpasses FVP in performance.

### v3

@@ 1 Introduction
In general, there are two ways of exploiting SAM in the SFDA scenario. One is to finetune SAM (usually by injecting adapters), with the algorithms in, e.g., SAMed [18], MA-SAM [19] and Med-SA [20]. However, in SFDA where the ground truth is unavailable, the performance of finetuning SAM highly depends on the quality of pseudo-labels. Bad pseudo-labels cannot yield satisfactory results (see Section 5.3.3). Another manner of leveraging SAM for SFDA is generating superior segmentation masks with the counterparts of SAM in the medical domain, e.g., MedSAM [21] and SAM-Med3D [22], to improve SFDA results, where the bottleneck is auto-generated valid prompts. Among the prompting modes of the medical domain SAM, the bounding box prompt can clearly specify the segmentation target without ambiguity and yields better segmentation accuracy compared to the point prompt[21, 22]. Thus, we aim to offer bounding box prompts to SAM, such as MedSAM[21], for enhancing domain adaptation. Existing auto-prompting methods require supervised fine-tuning of SAM, which falls outside the scope of the SFDA paradigm[23, 24]. Naively using the predictions of the model resulting from existing SFDA methods to obtain box prompts for SAM shows limited performance gains (see the rows “ProtoContra w/ MedSAM” in Table 123). Therefore, a better method of automatically providing the bounding box prompt needs to be designed. This paper focuses on using MedSAM, a renowned medical domain SAM with box prompt.
@@ 4.1 Preliminaries
Denote the source domain by $\mathcal{D}^{s}=(\mathcal{X}^{s},\mathcal{Y}^{s})$ and target domain by $\mathcal{D}^{t}=(\mathcal{X}^{t},\mathcal{Y}^{t})$. In SFDA, a source model $F^{s}:\mathcal{X}^{s}\rightarrow\mathcal{Y}^{s}$ is trained with a source domain dataset $\{x^{s}_{i},y^{s}_{i}\}_{i=1}^{n_{s}}$, where $(x^{s}_{i},y^{s}_{i})\in(\mathcal{X}^{s},\mathcal{Y}^{s})$. In this paper, we consider the medical image segmentation task, so $F^{s}$ is usually trained with a supervision loss like cross-entropy loss or Dice loss. $F^{s}$ consists of a feature extractor $G^{s}$ and a classifier $H^{s}$, i.e., $F^{s}(x)=H^{s}(G^{s}(x))$. $G^{s}$ maps each pixel $i\!\in\!\{1,\!\cdots\!,h\!\times\!w\}$ in an image to the feature $f_{i}\!\in\!\mathbb{R}^{d_{F}}$, and $H^{s}$ projects pixel feature into the semantic label space with $K$ classes. Here $h\times w$ is the image size and $d_{F}$ is the dimension of the extracted feature in model $F$. The source model is provided to the target client side. Also, an unlabeled dataset $\{x^{t}_{i}\}_{i=1}^{n_{t}}$ from the target domain $\mathcal{D}_{t}$ is given, where $x^{t}_{i}\in\mathcal{X}^{t}$. SFDA aims to learn a target model $F^{t}:\mathcal{X}^{t}\rightarrow\mathcal{Y}^{t}$ with only the source model $F^{s}$ and the target unannotated dataset $\{x^{t}_{i}\}_{i=1}^{n_{t}}$.
@@ 5.2 Comparison with State-of-the-Art Methods
To validate the effectiveness of our proposed DFG, we compared it with five recent state-of-the-art (SOTA) SFDA for medical image segmentation methods: DPL [10], AdaMI [15], UPL [12], FVP [14], and ProtoContra [16]. DPL [10] denoised pseudo-labels via pixel uncertainty estimation and prototype estimation. AdaMI [15] minimizes the entropy of model predictions with a class-ratio regularizer. UPL [12] obtain pseudo labels and their uncertainty via diversified predictions. FVP [14] learns a visual prompt parameterized with some low-frequency parameters in the frequency space. ProtoContra [16] aligns target features with class prototypes and uses contrastive learning to train the pixels with unreliable predictions. For a fair comparison, we reimplement DPL, AdaMI, UPL, and ProtoContra based on their papers and released codes. We utilize the same segmentation network, data augmentations, and type of optimizers. The batch sizes are tuned to choose the optimal values, and all the other hyperparameters are unchanged. For FVP, we directly reference the results from their paper as the codes are not released.
@@ 5.3.3 Comparison with finetuning Segment Anything Model
Figure 7: Examples of failure cases of box prompt search. From left to right: Target images; Ground truth; Ground truth box prompt (red box), and box prompt search procedure (blue boxes), with boxes changing from small to large (For clarity, only the first, the last, and part of the other prompts are drawn); The number of pixels of changed MedSAM predictions when the box prompt is switched from the last to the current, with the horizontal axis being the index of the current prompt.
@@ 5.3.6 Analysis of efficiency and parameters
In terms of the number of parameters, DFG freezes MedSAM and trains only the target model without introducing additional trainable networks for adaptation. Namely, the trainable parameters are limited to the target segmentation network, which are the same as or fewer than previous SFDA methods, with the exception of FVP, which solely trains a visual prompt. Despite this, our method surpasses FVP in performance.